OpenStudy (anonymous):

Find the function given that it satisfies f''(x)=36x^2+24x and its graph has a horizontal tangent line at the point (0,1)

OpenStudy (anonymous):

first you must integrate F"(x) to get F'(x) F'(x) = \[\int\limits (36x ^{2}+24x) dx = 12x ^{3}+ 12x ^{2} +c\] because graph has horizontal tangent line at (0,1) so the value of gradient is 0 Substitute to F'(x) to find c 0 = 12(0)^3 + 12(0)^2 + c --> c = 0

OpenStudy (anonymous):

then you integrate F'(x) to get F(x) F(x) = \[\int\limits (12x ^{3}+12x ^{2}) dx = 3x ^{4}+4x ^{3}+ c\] Then substitute x = 0, F(x) = 1 1 = 3(0)^4 + 4(0)^3 + c --> c = 1 Hence, F(x) = \[3x ^{4}+4x ^{3}+1\]